Design a buffer that has a pH of 4.84 using one of the weak acid/conjugate base systems shown below.

Weak Acid Conjugate Base Ka pKa
HC2O4- C2O42- 6.4 × 10-5 4.19
H2PO4- HPO42- 6.2 × 10-8 7.21
HCO3- CO32- 4.8 × 10-11 10.32



How many grams of the potassium salt of the weak acid must be combined with how many grams of the potassium salt of its conjugate base, to produce 1.00 L of a buffer that is 1.00 M in the weak base?

grams potassium salt of weak acid =

grams potassium salt of conjugate base =
Answer #1

Solution :

The buffer is prepared by HC2O4- and C2O4^2- by using the Hendersen - Hasselbalch equation,

pH = pka + log [C2O4^2-] / [HC2O4-]

4.84 = 4.19 + log [C2O4^2-] /[HC2O4-]

log [C2O4^2-] /[HC2O4-] = 0.65

[C2O4^2-] /[HC2O4-] = antilog 0.65 = 4.47

Thus, for the pH of 4.84 of 1M, 1L buffer preparation,

[C2O4-] = 4.67 x [HC2O4-]

>> Mass of potassium salt of weak acid (HC2O4-)

= 1 mol x molar mass = 1 mol x 128.124 g mol-1 = 128.12 g

>> Mass of grams of potassium salt of conjugate base (C2O4^2-) = 4.67 mol x molar mass

= 4.67 mol x 166.22 g mol-1 = 776.24 g